OpenStudy (anonymous):

[9.02] Solve x2 − 8x − 20 = 0. x = 4, x = −5 x = −4, x = 5 x = 2, x = −10 x = −2, x = 10

OpenStudy (anonymous):

Solution of X^2-8x-20=0 is.... (x-10)(x+2) so answers x=10, -2 is it right

OpenStudy (anonymous):

Yes you are correct.

OpenStudy (anonymous):

If you always want to double check you can try using the quadratic formula. It's longer doesnt hurt to try.

OpenStudy (anonymous):

so its D right

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Dont forget to give me thumbs up!